OpenStudy (anonymous):

Calculate the area of triangle QRS with altitude ST, given Q (0, 5), R (−5, 0), S (−3, 4), and T (−2, 3).

OpenStudy (anonymous):

A- 6.2 square units B- 7 square units C- 5.9 square units D- 5 square units

OpenStudy (ajprincess):

r u familiar with the distance formula?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

x2−x1)2+(y2−y1)2

OpenStudy (anonymous):

@ml1042903

OpenStudy (anonymous):

@Mertsj

OpenStudy (ajprincess):

using that find the lenghts of QR and ST. Area=(1/2)*QR*ST

OpenStudy (anonymous):

Don't know sorry! :( Try @genius12

OpenStudy (ajprincess):

@britt5 did u find the lengths of QR and ST

OpenStudy (anonymous):

i am working on it

OpenStudy (ajprincess):

oh k. carry on.:)

OpenStudy (anonymous):

2

OpenStudy (ajprincess):

what's 2?

OpenStudy (anonymous):

that what i got

OpenStudy (ajprincess):

for what did u get 2? for the length of QR?

OpenStudy (ajprincess):

@britt5

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

it is nt right:( QR=sqrt((-5-0)^2+(0-5)^2) =sqrt(25+25) =sqrt(50) =5*sqrt(2)

OpenStudy (anonymous):

i knew i looked wrong

OpenStudy (anonymous):

@ajprincess so the answer is D- 5 square units

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

Thanks can you help me on some more

OpenStudy (ajprincess):

ya post them:)

OpenStudy (anonymous):

A segment with endpoints I (5, 2) and J (9, 10) is divided by a point K such that IK and IJ form a 2:3 ratio. Find the y value for K.

OpenStudy (anonymous):

(A) 4.6 (B) 5.4 (C) 4.8 (D) 5.2

OpenStudy (anonymous):

@ajprincess i think it is C 4.8

OpenStudy (anonymous):

Write an equation of a line perpendicular to y = 1 over 2x − 5 in slope-intercept form that passes through the point (2, −7).

OpenStudy (anonymous):

(A) y= 1 over 2x − 7 (B) y= 2x + 2 (C) y= −2x − 5 (D) y= −2x − 3

OpenStudy (anonymous):

@Gabylovesyou

OpenStudy (ajprincess):

for the last question, the product of slopes of perpendicular lines is equal to -1. the slope of the given line is (1/2). let m be the slope of the required line. so m*(1/2)=-1 m=-2 slope intercept form of equation is y=mx+c y=-2x+c plugging the given point will give the value of c -7=-2*2+c -7=-4+c -7+4=c -3=c

OpenStudy (anonymous):

@ajprincess So do yo think the other one is C 4.8 is the last one (D) y= −2x − 3

OpenStudy (ajprincess):

I am working on the second one. but last one is option D

OpenStudy (anonymous):

Ok thanks on the second one i got 3.6 it is equivalent to 4.8

OpenStudy (anonymous):

Line IJ has an equation of a line y = 3x − 8, and line KL has an equation of a line y = 3x + 4. These two lines are A) parallel because the slopes of the lines are the same B) perpendicular because the product of the slopes is −1 C.) both representations of the same line D) neither parallel nor perpendicular

OpenStudy (anonymous):

C.) both representations of the same line

OpenStudy (anonymous):

Write an equation of a line parallel to line EF below in slope-intercept form that passes through the point (2, 6). A) y = 2x + 3 B) y = 3x + 6 C.) y = negative two over 3x + twenty eight over three D) y = − two over threex + 6

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

for the second one 3.6 or 4.8 is not right:( can u tell me how u workd it out?

OpenStudy (ajprincess):

may i knw y u say that both represent the same line

OpenStudy (anonymous):

For the second one i used a ratio calculator

OpenStudy (anonymous):

) perpendicular because the product of the slopes is −1

OpenStudy (ajprincess):

hmm. r u sure the product of the slopes is -1?

OpenStudy (anonymous):

the second one would it be 4.6

OpenStudy (ajprincess):

nope.

OpenStudy (anonymous):

B 5.4

OpenStudy (ajprincess):

say the total distance between I and J be 5d. 5d=sqrt((9-5)^2+(10-2)^2) =sqrt(4^2+8^2) =sqrt(16+64) =sqrt(80) =4*sqrt5 then d=(4*sqrt5)/5 the distance between I and K according to the given ratio 2d=(8*sqrt5)/5=sqrt((x-5)^2+(y-2)^2) the distance between K and J will be 3d=(12*sqrt5)/5=sqrt((9-x)^2+(10-y)^2) solve these two equations to get y.

OpenStudy (anonymous):

5.2

OpenStudy (ajprincess):

yup

OpenStudy (anonymous):

For the line segment whose endpoints are R (1, 2) and S (6, 7), find the y value for the point located 3 over 4 the distance from R to S. A) 4.25 B) 4.75 C.) 5.25 D) 5.75

OpenStudy (ajprincess):

before solving this one have u figured out the answer for 4th ques?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i only have that one and this one left

OpenStudy (ajprincess):

find the slopes of the two equations and see if they are equal or the product is -1.

OpenStudy (anonymous):

and the question with the picture i still have

OpenStudy (ajprincess):

I am sry I'v gotta go nw. It's past midnight here.

OpenStudy (anonymous):

y = 3x + 4 m=-3 y = 3x − 8 m=-3 A)slope are the same

OpenStudy (ajprincess):

m=3 nt -3 and yes slopes are same. If the slopes are same they are parallel.

OpenStudy (anonymous):

Ok good night thank so much

OpenStudy (anonymous):

ill guess on the other ones